Suppose the population over 15 is 20 million, the labour force Participation rate is 60%, and the unemployment rate is 5%. If the labour force grows by 3% and employment grows by 2%, what does the unemployment rate become?
1
Expert's answer
2019-06-05T09:23:23-0400

Current situation:

Labor force: 0.6*20=12 million

Number of unemployed persons: 0.05*12=0.6 million

Number of employed persons: 12-0.06=11.4 million

Supposed situation:

Labor force: 12*1.03=12.36

Number of employed persons: 11.4*1.02=11.628 million

Number of employed persons: 12.36-11.628=0.732 million

Unemployment rate:  0.732/12.36=5.9%